OpenStudy (anonymous):

Find the equation in standard form which passes through (-2, -7) and (1, -1)

OpenStudy (anonymous):

(-7-(-1))/(-2-1)=-6/-3=2 slope y=2x+b -1=2*1+b -1=2+b -3=b y=2x-3